Determine the elements of the following sequence

1. A=n^2-3n+3 where: 4<n<8

2. B = n (-1) ^n where: 2<n<4


1
Expert's answer
2021-03-30T06:45:33-0400

"A(5)=13, A(6)=21, A(7)=31. B(3)=-3." However I think the problem is not very correct. Actually may be it was meant, "A=\Sigma( n^2-3n+3)." Hence "A= \Sigma n^2 -3\Sigma n+\Sigma3=\frac{n(n+1)(2n+1)}{6}-3\frac{n(n+1)}{2}+3n" "=\frac{n(n^2-3n+5)}{3}." And B is given by "\Sigma (-1)^nn." For this let the sum be upto "2n" terms. Then subtracting "\Sigma n" we get the sum as "-2(1+3+\cdots +2n-1)=-2""n^2" . Hence required sum is "-2n^2 =B-\frac{2n(2n+1)}{2}\Rightarrow B=n". If number of terms is "2n+1," then "B= n-(2n+1)=-n-1."